# Context, Stochasticity, and Meaning in Large Language Models

### v2

@@ Abstract
We analyze the resolution of polysemy through dynamic, context-aware vector embeddings, contrasting this with a quantum-inspired framework for semantic context. We then investigate the intentional role of stochasticity in the generative process, identifying it as the core mechanism for creative and varied outputs. Finally, we explore the compelling analogy between the model’s reliance on vector-space projections and the formalisms of quantum mechanics. These principles are presented within the complete training context—from pre-training to reinforcement learning—to provide a cohesive understanding of both the construction and the intrinsic behavior of these models.
@@ VI.2 An Alternative Framework: Quantum Contextuality
Figure 1: Hypergraph visualization of three intertwining contexts for the word ‘bank’. The central vector $\bm{v}(\text{bank})$ is a shared element common to three distinct contexts (orthonormal bases). Each context connects $\bm{v}(\text{run})$ to a different semantic association: physical activity ($\bm{v}(\text{exercise})$), management ($\bm{v}(\text{business})$), and flow ($\bm{v}(\text{river})$).
This quantum-like contextual model of word encoding bears great similarities to quantum logic. The two main entities entering the discussion are (i) the vectors as well as the (ii) orthogonal (orthonormal) bases (or contexts or blocks or maximal operators) those vectors could be in. Think of, say, the Cartesian standard basis $\{\bm{e}_{1},\bm{e}_{2},\bm{e}_{3}\}$ with $\bm{e}_{1}=(1,0,0)^{\intercal}$, $\bm{e}_{2}=(0,1,0)^{\intercal}$, and $\bm{e}_{3}=(0,0,1)^{\intercal}$ as well as the orthogonal basis $\{\bm{f}_{1},\bm{f}_{2},\bm{f}_{3}\}$ with $\bm{f}_{1}=(1/\sqrt{2},-1/\sqrt{2},0)^{\intercal}$, $\bm{f}_{2}=(1/\sqrt{2},1/\sqrt{2},0)^{\intercal}$, and $\bm{f}_{3}=(0,0,1)^{\intercal}$, obtained by rotating the first and second vectors by an angle $\pi/4$ clockwise around the z-axis, which is aligned with $\bm{e}_{3}$. In this construction both bases share the same vector $\bm{e}_{3}=\bm{f}_{3}$, but the context or maximal observables involved are very different: in one case the maximal observable, in its spectral decomposition with arbitrary real eigenvalues $\lambda_{i}$, is:
@@ IX Comparison to Quantum Mechanics
Could we live as embedded observers [19] in an LLM? Large Language Models (LLMs) and quantum mechanics (QM) both formulate their fundamental objects as vectors and make decisive use of inner products and choices of basis. In QM, physical states are rays in a complex Hilbert space, with measurement statistics governed by projections onto an orthonormal basis (the “measurement context”). In LLMs, symbols and features are embedded as real-valued vectors, where attention and output generation are dominated by dot products against context-dependent representations. This shared linear-algebraic foundation invites a careful analogy and raises a speculative but concrete question: in what sense could an observer exist inside an LLM?
@@ IX.1 Vectors, Inner Products, and Contexts
In contrast, a Transformer computes layer-wise updates for its hidden state representations. A complete Transformer block applies two main sub-layers, each followed by a residual connection and layer normalization. The act of sampling a token is akin to a measurement that both reads out and updates the system’s state. The process is stochastic and irreversible due to sampling and normalization. Yet, it again features context-dependent projection: the attention mechanism pools information based on dot-product similarity, and the output layer projects the final hidden state onto basis-like directions defined by the rows of $\bm{W}_{\text{out}}$.
@@ IX.5 Orthogonal Bases and Functional Decomposition
Both QM and LLMs rely on decompositions into approximately orthogonal components to render predictions tractable. In QM, orthonormal bases diagonalize commuting observables. In LLMs, learned representations can be analyzed in bases where features are sparse and nearly orthogonal. The attention mechanism often acts as a soft projector onto subspaces spanned by key vectors most aligned with the current query. From an internal perspective, an observer that leverages such a basis could better ensure its persistence against interference from unrelated computations. Orthogonality provides robustness, much as pointer states do in decoherence theory.
@@ IX.6 Analogies, Disanalogies, and Conclusion
X Summary and Outlook

### v3

# From Polysemy to Quantum Attention: Context, Randomness, and Hilbert Space Perspectives on Large Language Models

@@ Abstract
We trace a conceptual path from linguistic ambiguity to quantum-inspired architectures in the study of Large Language Models (LLMs). First, we examine polysemy and show how modern transformer networks resolve multiple word meanings through dynamic, context-dependent embeddings, contrasting this with an alternative framework rooted in quantum contextuality. Next, we consider stochasticity not as a flaw but as a central design principle: randomness in sampling provides the mechanism by which LLMs generate diverse, creative outputs and explore their own representational landscape. Finally, we introduce the notion of quantum attention, a speculative extension of the classical attention mechanism that constrains its transformations to be unitary. This shift reframes attention as a reversible rotation in Hilbert space rather than a dissipative mixture, aligning the processing of meaning in language models with fundamental structures of quantum theory.
@@ VI.2 An Alternative Framework: Quantum Contextuality
Figure 1: Hypergraph visualization of three intertwining contexts for the word ‘bank’. The central vector $\bm{v}(\text{bank})$ is a shared element common to three distinct contexts (orthonormal bases). Each context connects $\bm{v}(\text{run})$ to a different semantic association: physical activity $\bm{v}(\text{exercise})$, management $\bm{v}(\text{business})$, and flow $\bm{v}(\text{river})$.
This quantum-like contextual model of word encoding bears great similarities to quantum logic. The two main entities entering the discussion are (i) the vectors as well as the (ii) orthogonal (orthonormal) bases (or contexts or blocks or maximal operators) those vectors could be in. Think of, say, the Cartesian standard basis $\{\bm{e}_{1},\bm{e}_{2},\bm{e}_{3}\}$ with $\bm{e}_{1}=(1,0,0)^{\intercal}$, $\bm{e}_{2}=(0,1,0)^{\intercal}$, and $\bm{e}_{3}=(0,0,1)^{\intercal}$ as well as the orthonormal basis $\{\bm{f}_{1},\bm{f}_{2},\bm{f}_{3}\}$ with $\bm{f}_{1}=(1/\sqrt{2})\,(1,-1,0)^{\intercal}$, $\bm{f}_{2}=(1/\sqrt{2})\,(1,1,0)^{\intercal}$, and $\bm{f}_{3}=(0,0,1)^{\intercal}$, obtained by rotating the first and second vectors by an angle $\pi/4$ clockwise around the z-axis, which is aligned with $\bm{e}_{3}$. In this construction both bases share the same vector $\bm{e}_{3}=\bm{f}_{3}$, but the context or maximal observables involved are very different: in one case the maximal observable, in its spectral decomposition with arbitrary real eigenvalues $\lambda_{i}$, is:
@@ VII The Origin of Stochasticity in responses
Even if the temperature is set to zero, a residual stochasticity persists due to the parallelization and batch processing inherent in large-scale inference. At a temperature of 0, the source of nondeterminism is the absence of batch invariance in highly parallel kernel use: Any fluctuating load determines the variations of the batch size into which any given request is grouped and parallelized. This variation in batch size causes high-performance kernels for operations like matrix multiplication, normalization, and attention to alter their internal reduction strategy to maintain efficiency. At a fundamental level, a reduction is a sequence of floating-point additions; the reduction “strategy” specifies how these additions are partitioned and executed across parallel compute units. Because floating-point addition is lossy for two floating-point numbers with different exponents, and thus non-associative—that is, $(a+b)+c$ is not guaranteed to be bit-wise identical to a+(b+c)a+(b+c)—any change in the reduction order can produce a numerically different result. Consequently, the final output of an individual request becomes dependent on the (for the user seemingly) nondeterministic batching process [10].
@@ IX Comparison to Quantum Mechanics
Could we live as embedded observers [20] in an LLM? LLMs and quantum mechanics (QM) both formulate their fundamental objects as vectors and make decisive use of inner products and choices of basis. In QM, physical states are rays in a complex Hilbert space, with measurement statistics governed by projections onto an orthonormal basis (the “measurement context”). In LLMs, symbols and features are embedded as real-valued vectors, where attention and output generation are dominated by dot products against context-dependent representations. This shared linear-algebraic foundation invites a careful analogy and raises a speculative but concrete question: in what sense could an observer exist inside an LLM?
@@ IX.1 Vectors, Inner Products, and Contexts
In contrast, a transformer computes layer-wise updates for its hidden state representations [21, 22]. A complete Transformer block applies two main sub-layers, each followed by a residual connection and layer normalization. The act of sampling a token is akin to a measurement that both reads out and updates the system’s state. The process is stochastic and irreversible due to sampling and normalization. Yet, it again features context-dependent projection: the attention mechanism pools information based on dot-product similarity, and the output layer projects the final hidden state onto basis-like directions defined by the rows of $\bm{W}_{\text{out}}$.
@@ IX.5 Orthogonal Bases and Functional Decomposition
Both QM and LLMs rely on decompositions into approximately orthogonal components [23] to render predictions tractable. In QM, orthonormal bases diagonalize commuting observables. In LLMs, learned representations can be analyzed in bases where features are sparse and nearly orthogonal. The attention mechanism often acts as a soft projector onto subspaces spanned by key vectors most aligned with the current query. From an internal perspective, an observer that leverages such a basis could better ensure its persistence against interference from unrelated computations. Orthogonality provides robustness, much as pointer states do in decoherence theory.
@@ IX.6 Analogies, Disanalogies, and Conclusion
X Quantum Attention
The architectures of modern Large Language Models have inherited much from classical attention mechanisms, yet attempts to “quantize” these ideas have often been limited to almost literal translations of the classical recipe into the language of Hilbert spaces [22]. Such approaches, while intuitive, miss the more radical point: quantum evolution admits only two fundamental capacities, identified by von Neumann as (i) irreversible, stochastic state reduction (Process 1), and (ii) reversible, deterministic unitary evolution (Process 2). A viable notion of quantum attention must therefore be constructed wholly within these bounds.
X.1 Goals of Quantum Attention
The principal objective of attention in the classical setting is to guide next-token prediction through dynamically weighted embeddings. Translated into a genuinely quantum language, this entails:
  1. 1.
    Encoding tokens as quantum states $\bm{\psi}$ in a high-dimensional Hilbert space, thereby embedding them into the “meaning-space” of the model.
  2. 2.
    Evolving these states by a suitably chosen unitary transformation $\bm{U}$, producing a candidate output state
    $\bm{\phi}=\bm{U}\bm{\psi}.$ (7)
  3. 3.
    Performing a terminal measurement (Process 1), which yields a single token according to the Born rule, with probabilities proportional to $|\bm{\phi}^{\dagger}\bm{\psi}|^{2}$ for pure states.
In this scheme, nonlinearities and irreversibility are absent everywhere except at the initial state preparation (if stochastic) and at the final, irreducibly random measurement. Everything in between must be faithfully unitary.
X.2 Constraints and Learning Mechanisms
Because all allowed intermediate transformations fall within class (ii)—unitary processes—training and back-propagation must be reconceived as unitary rotations in Hilbert space. This sharply contrasts with the gradient-descent-based adjustments of non-linear maps in classical neural networks. Concretely, learnable transformations are constrained to be:
  • In the general case, arbitrary unitaries acting on the relevant Hilbert space, implemented, e.g., by parametrized quantum circuits.
  • In restricted scenarios, orthogonal transformations over real-valued state spaces, or, in an extreme discrete approximation, mere permutations of basis states.
Thus, what counts as “attention weights” must be replaced not by scalar coefficients, but by controlled rotations or permutations of the embedding subspace.
X.3 Comparison with Classical Attention
Classical attention is a fundamentally nonlinear mechanism: the query–key compatibility scores undergo a softmax normalization and act as dynamic coefficients for a weighted average of value embeddings. Quantum attention, by contrast, cannot implement such probabilistic normalization mid-process. The only stochastic step is the final measurement. Hence, quantum attention is more austere but also more physically faithful: weights are rotations, superpositions evolve unitarily, and probabilities emerge only at the very end.
One may summarize the distinction with a slogan: classical attention mixes; quantum attention rotates. The former continuously interpolates between embeddings using nonlinear weightings, while the latter conserves norm and mixes information strictly through unitarity. Whether this stricter yet cleaner formalism can achieve the flexibility and expressivity that classical attention enjoys remains both a technical and philosophical open question.
X.4 Implications
The implications are sobering and exciting at once. On the sobering side, many of the workhorse techniques of classical deep learning, such as gradient backpropagation through non-linear activations, are unavailable in their standard form. On the exciting side, however, quantum attention reorients learning toward symmetry operations, reversible dynamics, and combinatorial richness—offering the prospect of models that are less fragile, more interpretable, and governed by deep conservation laws. If classical transformers are powered by the calculus of weighted averages, their quantum cousins would be powered by the geometry of Hilbert space rotations. “Paying attention,” in such a universe, would mean aligning one’s basis just right, before the final measurement tosses its fundamental dice.
XI Summary and Outlook